OpenStudy (anonymous):

how to rewrite 1/3s-8=t for s

OpenStudy (anonymous):

the first step would be to add the 8 to the other side. s0 1/3s = t+8

OpenStudy (anonymous):

do u know what the next step is to get s by itself ?

OpenStudy (anonymous):

you step the problem up wrong because the problem was 1/3s-8=t for s

OpenStudy (anonymous):

um can u draw it or write it in the equation editor

OpenStudy (anonymous):



OpenStudy (anonymous):